OpenStudy (anonymous):

HELP PLEASE!!! Think of an instance in your everyday life where you could use the help of a system of linear equations.

OpenStudy (anonymous):

what is question

OpenStudy (anonymous):

show me your question

OpenStudy (solomonzelman):

Easy !! INSTANCE: I am an owner of a store that sells computers. there are two factories/companies in town that produce them, sell (to stores) and deliver them in big numbers. Company 1 charges 2000 dollars for the delivery of all computers together, and for each computer (the cost to buy is) 200 dollars. Company 2 charges 1000 dollars for the delivery of all computers together, and for each computer (the cost to buy is) 300 dollars. So I, the owner of a store want to buy a certain number of computers for a store. How would it be cheaper to buy from company 1 or company 2? SOLUTION: write linear equation for each company. Company 1: y=200x+2000 Company 2: y=300x+1000 (where IN BOTH EQUATIONS x is the number of computers, y-intercept is the delivery fee, and y is the total cost. if you set them equal to each other you will find the number computers (x) that would cost the same in total. 200x+2000=300x+1000 solve for x, STEP 1: 200x + 2000 = 300x + 1000 -200x -200x you get 2000=100x+1000 STEP 2: 2000 = 100x + 1000 -1000 -1000 you get 1000=100x STEP 3: (final step) 1000 = 100x ÷10 ÷10 x=10 SO, when you buy 10 computer you will have to pay the same price in both companies. From 11 computers and more a cheaper deal (that costs less for the buyer/store-owner ) company 1 is better (better meaning cheaper) (Prove) Company 1: y=200x+2000 x=11 , 200(11)+2,000=2,200+2,000=3,800 Company 2: y=300x+1000 x=11 , 300(11)+1,000=3,300+1,000=4,300 From 9 computers and less company 2 is better. (better meaning cheaper) (Prove) Company 1: y=200x+2000 x=9 , 200(9)+2000=1800+2000=3800 Company 2: y=300x+1000 x=9 , 300(9)+1000=2700+1000=3700 NEED MORE HELP ?